OpenStudy (anonymous):

Let n be a positive whole number. i) Prove that nx2nCn = (n+1)x2nCn-1 ii) Use that fact that 2nCn - 2nCn-1 is a whole number to prove that 1/n+1x2nCn is a whole number Can you advise on method please. Thanks.

OpenStudy (experimentx):

just write it up, it's not difficult to prove it ... although I believe you rather want combinatorial argument why?

OpenStudy (experimentx):

for i) think about the ways forming a committee of n members from total of 2n members and selecting a caption from n members. of think about selecting n-1 committee members first and a captain from remaining members.

OpenStudy (experimentx):

ii) \[ \binom{2n}{n}-\binom{2n}{n-1} = \frac{1}{n+1} \left ( (n+1)\binom{2n}{n}-(n+1)\binom{2n}{n-1}\right ) \\ = \frac{1}{n+1} \left ( (n+1)\binom{2n}{n}-(n)\binom{2n}{n}\right ) = \frac{1}{n+1} \binom{2n}{n}\]

OpenStudy (experimentx):

I believe it is called Catlan number. http://en.wikipedia.org/wiki/Catalan_number